OpenStudy (perl):

A boat travels at a relative constant speed compared to the river of 15 mph upriver for 10 miles and back for 10 miles. The trip takes 1.5 hours. What is the speed of the current

OpenStudy (anonymous):

i don't really get the question. so the boat goes up the river 10 miles, then downstream the river 10 miles, takes 1.5 hours. the boat's speed > river's current's speed 15 mph?

OpenStudy (perl):

well ,

OpenStudy (perl):

the speed of the boat is 15 mph , i assume

OpenStudy (perl):

i assume the boat speed is 15 mph , relative to the water? or relative to

OpenStudy (perl):

yes the boat maintains a speed of 15 mph relative to the water

OpenStudy (perl):

let x=speed of current (relative to land) speed of boat downstream=15+x (relative to land) speed of boat upstream=15-x (relative to land)

OpenStudy (perl):

so this is like saying, the boat's speedometer is a constant 15 mph

OpenStudy (perl):

we have 10 = (15+x)*t1 , 10 = (15-x)*t2 10/(15+x) = t1, 10/(15-x) = t2 t1 + t2 = 1.5 10/(15+x) + 10/(15-x) =1.5